The ratio of boys to girls in a school is 5:4. If there are 500 girls, how many boys are there in the school?
Answer by Maths68(1474)

Let
Boys = b
Girls = g
Girls = g = 500
Boys = b = 500-x
b:g=5:4
b/g=5/4
4b=5g
Put the values of g and b
4(500-x)=5*500 ............(1)
2000-4x=2500
-4x=2500-2000
-4x=500
-4x/-4=500/-4
x=-125
Boys = 500-(-125)
Boys = 500+125
Boys = 625

Check
=======
Boys/girls=625/500
Boys/girls=125/100
Boys/girls=5/4
Boys:girls=5:4